OpenStudy (anonymous):

Can someone explain to me, in lamens terms (Pretend I am Extremely Dumb) how to calculate the square root of a large number, with four or five digits, without a calculator? Say The Sqrt(50989)

OpenStudy (anonymous):

*Layman's Terms then, if you please.

OpenStudy (anonymous):

Still don't get it, at all -_-

OpenStudy (aum):

50989 100 is a perfect square and so you can start from the RIGHT and skip every two digits and you will be left with 5. Estimate sqrt(5). The sqrt(50989) will be in the vicinity of 2*10*10

OpenStudy (anonymous):

Why would you start with the idea that 100 is a perfect square?

OpenStudy (anonymous):

Interesting that you'd pick such a large prime number...

OpenStudy (anonymous):

Let's say I did that on purpose.

OpenStudy (aum):

Let us say the number is 40,000 40,000 = 4 * 100 * 100 square root = 2 * 10 * 10 = 200

OpenStudy (anonymous):

Alright. 40, 000 = 4 * 100 * 100, I'm with you there.

OpenStudy (aum):

50989 is approx 5 * 100 * 100 so we get the ballpark estimate of sqrt(5) * 10 * 10 approx 200. Now we can refine it by trial and error.

OpenStudy (anonymous):

You know what, this is so hard that I'm going to stick with it until I can do these in my head. By hand first.

OpenStudy (anonymous):

As in sqrt(5 * 100 * 100) Ok, cool, I get that

OpenStudy (anonymous):

I don't want trial and error, I want an exact method! Lol!

OpenStudy (anonymous):

Another way, assuming you were given the square root of a number close to \(\sqrt{50989}\approx 225.807\), such as \(\sqrt{50625}=225\), you can use the linear approximation technique to estimate the first square root. Take the function \(f(x)=\sqrt x\), then \[f(50989)\approx f(50625)+f'(50625)(50989-50625)\]

OpenStudy (anonymous):

You just mind grenaded me Siths

OpenStudy (anonymous):

Gimme a min to look at this

OpenStudy (anonymous):

If you haven't seen anything about the linear approx. yet: http://tutorial.math.lamar.edu/Classes/CalcI/LinearApproximations.aspx

OpenStudy (anonymous):

I sure haven't. Let me do that tutorial. I'll be right back.

OpenStudy (aum):

You can also look up Newton-Raphson or just Newton method. But it is a successive approximation method.

OpenStudy (anonymous):

Wait, isn't the f'(x) when talking about \[f(x) = \sqrt[3]{x} \] equal to x^3?

OpenStudy (anonymous):

I haven't taken any calculus, is that an issue, or can I still get this?

OpenStudy (aum):

It is cube root

OpenStudy (anonymous):

Well thanks for the help anyways. I'll figure it out.

OpenStudy (aum):

You will have to know how to find the derivative of a function which is taught in calculus. But if all you are interested in is the square root function you can just memorize the derivative.

OpenStudy (anonymous):

Nah, I'm interested in learning Calculus much more than I am about learning this square root by hand operation. I need to be in the equivalent of Calculus 3 in less than 6 months. Then I have about 8 years of college ahead of me lol